OpenStudy (anonymous):

find a-b a=6i-6j , b=4i+2j options are: 12 -36 <24,-12> <10,-4> please help asap. will be highly appreciate. <3

OpenStudy (anonymous):

a-b=6i-6j-[4i+2j] =6i-6j-4i-2j=(6-4)i-(6+2)j=10i-8j i.e.<10,-8> :)

OpenStudy (anonymous):

Thanks a lot :)